OpenStudy (anonymous):

What is the 20th term of the arithmetic sequence 5, -1, -7, … ?

OpenStudy (cwrw238):

d = common difference = -1 - 5 = -6 nth term = a + (n-1)d

OpenStudy (anonymous):

n is 20 , right ?

OpenStudy (jiteshmeghwal9):

look carefully to the question,every term is adding -6 with it to gt the next term.

OpenStudy (cwrw238):

here a = 5 d = -6 and a = 20

OpenStudy (anonymous):

Yes..

OpenStudy (cwrw238):

sorry n = 20

OpenStudy (jiteshmeghwal9):

take this pattern to 20th term .u will probably gt ur answer

OpenStudy (anonymous):

So: \[t_{20} = 5 + (20-1)(-6)\]

OpenStudy (cwrw238):

yep

OpenStudy (anonymous):

-120 ?

OpenStudy (anonymous):

NO try it once more.. What is 19*-6) = ??

OpenStudy (anonymous):

19*(-6) = ???

OpenStudy (anonymous):

-114 . sorry i did 20x6-5

OpenStudy (anonymous):

-114 + 5 = ??

OpenStudy (anonymous):

-109

OpenStudy (cwrw238):

yea

OpenStudy (anonymous):

Yes absolutely..

OpenStudy (anonymous):

YAY! thanks :) there will probably be more questions like this to come ... :/ i'm awful at sequences & series :(

OpenStudy (anonymous):

Welcome dear..